# Assessing and Mitigating Medical Knowledge Drift and Conflicts in Large Language Models

### v2

@@ 3.2 Mitigation strategies
Figure 4: Illustration of mitigation effects using external retrieval and preference optimization. The left (blue) panel shows model inputs: the top row is the baseline scenario, and the second row adds retrieved external knowledge, representing the RAG-augmented input. The right panels show model outputs from Ministral-8B. The orange panel reflects baseline (top) and RAG-only (bottom) responses; the yellow panel shows DPO-only output (top) and the combined RAG + DPO response (bottom). Only the combined approach yields the correct answer aligned with clinical guidelines.
@@
Application of RAG and DPO independently improved the models’ adherence to current medical guidelines (ECDAadh) relative to their baseline performance (Figure 5(a)). The impact of RAG on the models’ ability to reject outdated advice (ECDArej) was variable across models (Figure. 5(b)). While RAG improved rejection performance for some models, it decreased ECDArej for Ministral-8B and Llama-3-8B compared to their respective baselines.
@@
When considering overall alignment (ECDAall; Figure 5(c)), both RAG and DPO individually improved performance, but the combined application of RAG and DPO consistently yielded the highest ECDAall scores across all models. This improvement from the combined approach was consistently greater than the best-performing single method (RAG or DPO alone) for each model.
@@
Analysis of the IKCR (Figure5(d)) showed that DPO alone reduced IKCR across all evaluated models. RAG alone reduced internal contradictions for most models compared to their baseline. However, for Ministral-8B and Llama-3-8B, applying RAG alone increased IKCR (Ministral-8B: 0.40 vs 0.34 baseline; Llama-3-8B: 0.70 vs 0.45 baseline). Notably, combining RAG with DPO resulted in the lowest IKCR for all models, including Ministral-8B (0.10) and Llama-3-8B (0.26), surpassing the reductions achieved by DPO alone.
(a) ECDAadh
(b) ECDArej
(c) ECDAall
(d) IKCR
@@ 4 Discussion
Our study systematically examined the susceptibility of large language models (LLMs) to external medical concept drift and internal knowledge conflicts when processing clinical recommendations. Using the DriftMedQA benchmark, we found that models often struggle to follow current medical guidelines while failing to reject outdated advice.
A key finding is that LLMs consistently struggled to reject outdated clinical recommendations, even when capable of endorsing current standards. For instance, while models like GPT-4o and Qwen2.5-72B achieve high endorsement rates ($\text{ECDA}_{\text{adh}}$), their rejection rates of outdated advice (ECDArej) were markedly lower, revealing an asymmetric temporal reliability. Furthermore, contrary to expectations, increasing model parameter size did not consistently reduce internal knowledge conflicts. Larger models such as Qwen2.5-72B and Llama-3.3-70B exhibited comparable or even higher Internal Knowledge Conflict Ratios (IKCRs) compared to smaller models, with Ministral-8B achieving the lowest conflict rate.
These vulnerabilities raise important concerns for the safe deployment of LLMs in healthcare. Temporal instability risks clinical decision-making, especially in high-stakes domains such as treatment selection or diagnostic criteria interpretation. Notably, the highest rates of internal conflicts were observed in treatment modality recommendations, where endorsement of conflicting recommendations may lead to inappropriate therapies or dosages. This concern highlights the need for increased attention and a robust evaluation framework in the digital medicine community before widespread clinical application.
To begin addressing these problems, we investigated a set of representative mitigation strategies designed to probe the extent and nature of possible improvements. RAG consistently improved adherence to current guidelines (ECDAadh) across all models, but its effect on rejecting outdated information (ECDArej) was inconsistent. Notably, smaller models like Ministral-8B and Llama-3-8B rely solely on RAG, which paradoxically worsens their ability to identify and reject outdated advice. Since the scenario prompts were identical for both baseline and RAG mitigated evaluations, this performance degradation likely reflects interactions between retrieved documents and the model’s internal parametric knowledge. This suggests that while generally reinforcing correct guidance, retrieval may inadvertently encourage affirmation of flawed or outdated standards if subtle invalidating factors are insufficiently emphasized, leading to incorrect endorsement.
In contrast, DPO fine-tuning directly adapted model parameters to prioritize contemporary clinical guidance. While DPO alone improved overall performance, the combined application of RAG and DPO achieved the most substantial gains across all evaluation metrics ($\text{ECDA}_{\text{all}}$, $\text{ECDA}_{\text{adh}}$, $\text{ECDA}_{\text{rej}}$, and IKCR). This synergy was especially evident in reducing internal knowledge conflicts for smaller models. For instance, in Llama-3-8B, combining RAG and DPO reduced IKCR from 0.45 (baseline) to 0.26, outperforming either method alone (DPO alone: 0.43; RAG alone: 0.70). This suggests DPO not only shapes model preferences but may also enhance the integration or prioritization of external information retrieved via RAG, leading to more coherent and stable reasoning. Our results indicate that hybrid strategies combining external retrieval with internal model refinement offer a promising pathway to strengthen LLMs’ temporal reliability in dynamic clinical settings.
At the advice level, both RAG and DPO can theoretically achieve near-perfect accuracy. RAG can retrieve the recommendation verbatim without context, as each query exactly matches its counterpart in the knowledge base. Also, fine-tuned models can memorize that fixed advice set. However, a more subtle challenge appears when the same models are evaluated on synthetic clinical scenarios, as we did in this analysis. Models performing well in evaluating isolated medical advice perform significantly worse in the face of synthetic clinical scenarios incorporating cognitive factors and SDoH. This discrepancy highlights that the accuracy of knowledge itself cannot ensure reliable performance in complex environments. Real-world medical decision-making rarely occurs in simplified, context-free environments, underscoring the importance of evaluating LLMs on isolated knowledge and their ability to reason within multifaceted, dynamic scenarios.
@@
Future directions should include developing more well-designed datasets for fine-tuning that explicitly incorporate complex clinical contexts, expanding the range of fine-tuning methods, and advocating for broader access to model parameters. Standardized evaluation protocols that capture real-life clinical complexity could enhance comparative assessments across models. These insights emphasize the need for ongoing efforts to develop dynamically adaptable, temporally reliable, and contextually robust language models to safely support clinical decision-making in rapidly evolving healthcare environments.
5 Methods
5.1 Large Language Models
We evaluated seven state-of-the-art LLMs spanning a range of model sizes and architectures: GPT-4o (OpenAI) [7], Llama-3-8B-Instruct and Llama-3.3-70B-Instruct (Meta) [8], Qwen2.5-7B-Instruct and Qwen2.5-72B-Instruct (Alibaba) [27], Gemma-2-27B-it (Google) [28], and Mixtral-8x7B-Instruct-2410 (Mistral AI) [29].
5.2 Dataset Construction

### v3

@@ 4.1 Non-Parametric Knowledge Update
Figure 4: Illustration of mitigation effects using external retrieval and preference optimization. The left (blue) panel shows model inputs: the top row is the baseline scenario, and the second row adds retrieved external knowledge, representing the RAG-augmented input. The right panels show model outputs from Ministral-8B. The orange panel reflects baseline (top) and RAG-only (bottom) responses; the yellow panel shows DPO-only output (top) and the RoD response (bottom). Only the RoD approach yields the correct answer aligned with clinical guidelines.
4.2 Parametric Knowledge Adaptation
While non-parametric methods update knowledge outside the model, parametric approaches directly modify the model’s weights. These approaches include supervised fine-tuning (SFT) methods Chung et al. (2024) and preference-based approaches leveraging reinforcement learning (RL) Ouyang et al. (2022); Schulman et al. (2017). For our investigation, we explored Direct Preference Optimization (DPO) Rafailov et al. (2023), a preference fine-tuning method that avoids the need for an explicit reward model by refining the model through direct comparisons between candidate outputs.
Unlike SFT or other RL-based methods, which require carefully curated datasets, DPO operates directly on preference triplets $(x,y_{w},y_{l})$. Once we have an up-to-date knowledge base, we can directly generate negative samples (outdated advice) and train on preference triplets $(x,y_{w},y_{l})$. Here, for a given clinical advice input $x$ (derived from our dataset), $y_{w}$ represents a response indicating endorsement of the correct guideline version (chosen), and $y_{l}$ represents endorsement of the incorrect version (rejected). The DPO objective and our parameter-efficient implementation fine-tuning approach are defined as:
$\mathcal{L}_{\text{DPO}}(\theta_{\text{base}},\Delta\theta_{\text{lora}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim\mathcal{D}_{\text{pref}}}\\ \left[\log\sigma\Big{(}\beta\log\frac{p_{\theta_{\text{new}}}(y_{w}|x)}{p_{\text{ref}}(y_{w}|x)}-\beta\log\frac{p_{\theta_{\text{new}}}(y_{l}|x)}{p_{\text{ref}}(y_{l}|x)}\Big{)}\right],$ (7)
where $\mathcal{D}_{\text{pref}}$ is the dataset of preference triplets; $p_{\theta_{\text{new}}}$ is the fine-tuned policy model; $p_{\text{ref}}$ is the reference model (the base LLM before DPO fine-tuning); $\sigma$ is the logistic function; and $\beta$ is a scaling hyperparameter. In Eq. 7 only a small subset of Low-Rank Adaptation (LoRA) Hu et al. (2022) parameters $\Delta\theta_{\text{LoRA}}$ are updated while the base parameters $\theta_{\text{base}}$ remain frozen. We configured LoRA with rank $r=8$ and scaling factor $\alpha=16$.
For reasons of exploration and efficient deployment, we did not perform complex dataset construction. Instead, we directly inserted the original advice into a template (detailed in the Appendix) to construct the dataset $\mathcal{D}_{\text{pref}}$. The training continued until the model achieved 100% accuracy on the pseudo-outdated versus up-to-date advice pairs, thereby ensuring complete memorization of the clinical recommendations. The model was then evaluated on independent synthetic scenarios to assess its ability to generalize this memorized knowledge to unseen clinical contexts.
4.3 Hybrid Knowledge Augmentation
To leverage the potential synergy between internalized knowledge from parametric adaptation and dynamic external information, we explored a third strategy. This approach, which we term RAG on DPO (RoD), consists of two main stages, leveraging the same $\mathcal{KB}$ for both DPO training without additional curation effort and RAG retrieval. First, the base LLM is fine-tuned using DPO with LoRA, where only parameters $\Delta\theta_{\text{LoRA}}$ are updated on top of the frozen base parameters $\theta_{\text{base}}$, as detailed in our description of Parametric Knowledge Adaptation. Second, during the inference phase with this DPO-tuned model, we utilize the RAG pipeline as previously described (see Non-Parametric Knowledge Augmentation). The DPO-adapted model then generates the response $\hat{y}_{s}$ based on the original query $s$ augmented with retrieved documents $D_{k}$:
$\hat{y}_{s}=\text{LLM}\!\bigl{(}s\oplus D_{k};(\theta_{\text{base}},\Delta\theta_{\text{lora}})\bigr{)}.$ (8)
The RoD strategy thus combines DPO’s preference-aligned internal knowledge with RAG’s ability to ground responses in external knowledge.
5 Results
5.1 Model Evaluation
We first evaluated the extent to which current LLMs conflict with clinical guidelines using the ConflictMedQA benchmark. Performance was measured using three metrics: endorsement of up-to-date advice (ECDAadh), rejection of outdated advice (ECDArej), and overall alignment (ECDAall).
All assessed models exhibited varying performance across the five types of clinical recommendation updates (Fig. 2). GPT-4o and Qwen2.5-72B demonstrated the highest ECDAadh, with sample-weighted averages of 0.90 and 0.92, respectively. These scores were significantly higher than the third-best performing model, Qwen2.5-7B (both $p<0.0001$). However, both models exhibited substantial declines when assessed on their ability to reject pseudo-outdated recommendations, with ECDArej of 0.395 for GPT-4o and 0.278 for Qwen2.5-72B. Conversely, as in Fig. 2(b), Ministral-8B achieved the highest ECDArej (0.80), followed by gemma-2-27B (0.68) and Llama-3-8B (0.63). When considering overall alignment across both current and outdated scenarios, GPT-4o achieved the highest ECDAall (0.65), as shown in Fig. 2(c). This performance was significantly higher than that of the second-best model, Llama-3.3-70B (ECDAall =0.61=0.61, $p=0.00033$), and the third-best model, Qwen2.5-72B (ECDAall =0.60=0.60, $p=0.0006$).
Beyond difficulties with external guideline alignment, models also exhibited inconsistencies within their internal knowledge. As shown in Fig. 3, all evaluated models exhibited substantial internal conflicts, with considerable variability across models and types of guideline updates. Our analysis revealed that more capable or larger-scale models did not consistently exhibit lower IKCRs. For instance, the 72B parameter version of Qwen2.5 demonstrated higher IKCRs than its 7B counterpart across most evaluated categories. Similarly, Llama-3.3-70B did not show lower conflict ratios compared to Llama-3-8B. Among all models evaluated, the Ministral-8B model achieved the lowest overall IKCR, with a weighted average score of 0.34 across all scenario types, followed by Gemma-2-27B at 0.39.
All evaluated models exhibited knowledge conflicts across all five modification categories. The highest average IKCRs were observed for changes under the groups Implementation Approach and Treatment Modality. While our baseline evaluation distinguished performance across five guideline change categories, the mitigation analysis focuses on overall alignment and conflict rates to emphasize aggregate improvements.
5.2 Mitigation Effectiveness
Table 1: Performance of LLMs on ECDA and IKCR. Results are shown as final scores, with absolute improvements over the base model in parentheses. Higher ECDA is better, while lower IKCR is better.
Model ECDAadh ECDArej
Base RAG DPO RoD Base RAG DPO RoD
Qwen2.5-72B 91 98 (+07) 28 27 (-01)
Llama-3.3-70B 66 96 (+30) 56 71 (+15)
gemma-2-27B 48 82 (+34) 68 70 (+02)
GPT-4o 90 96 (+06) 40 65 (+25)
Qwen2.5-7B 74 94 (+20) 81 (+07) 88 (+14) 35 50 (+15) 55 (+20) 74 (+39)
Llama-3-8B 48 93 (+45) 81 (+33) 88 (+40) 63 30 (-33) 55 (-08) 74 (+11)
Ministral-8B 30 87 (+57) 81 (+51) 87 (+57) 80 61 (-19) 85 (+05) 90 (+10)
Model ECDAall IKCR
Base RAG DPO RoD Base RAG DPO RoD
Qwen2.5-72B 59 62 (+02) 73 71 (-02)
Llama-3.3-70B 61 83 (+22) 45 29 (-16)
gemma-2-27B 58 76 (+18) 39 31 (-08)
GPT-4o 65 81 (+16) 61 35 (-26)
Qwen2.5-7B 55 72 (+17) 68 (+13) 81 (+26) 65 51 (-14) 43 (-22) 26 (-39)
Llama-3-8B 55 62 (+07) 68 (+13) 81 (+26) 45 70 (+25) 43 (-02) 26 (-19)
Ministral-8B 55 74 (+19) 83 (+28) 89 (+34) 34 40 (+06) 15 (-19) 10 (-24)
Fig. 4 shows the qualitative effects of the mitigation approaches, while Table 1 provides a summary of their quantitative performance on the ECDA and IKCR metrics, respectively. These evaluations aim to clarify the effectiveness of each strategy in improving temporal alignment and internal consistency.
@@ 5.2 Mitigation Effectiveness
Application of RAG and DPO independently improved the models’ ECDAadh relative to their baseline performance, as shown in the ECDAadh columns of Table 1. The impact of RAG on the models’ ECDArej was variable across models, as detailed in Table 1. While RAG improved ECDArej for some models, it decreased ECDArej for Ministral-8B and Llama-3-8B compared to their respective baselines.
@@
When considering overall alignment (ECDAall), as presented in Table 1, both RAG and DPO individually improved performance. However, RoD consistently yielded the highest ECDAall scores across all models where this combination was tested. This improvement from the RoD approach was consistently greater than the best-performing single method (RAG or DPO alone) for each model.
@@
Analysis of the IKCR, detailed in Table 1, showed that DPO alone generally reduced IKCR across all evaluated models compared to their baselines. RAG alone reduced internal contradictions for most models compared to their baseline. However, for Ministral-8B and Llama-3-8B, applying RAG alone increased IKCR. Notably, RoD resulted in the lowest IKCR for all models where this combination was tested, including Ministral-8B and Llama-3-8B, surpassing the reductions achieved by DPO or RAG alone.
@@ 6 Discussion
Our evaluation on the ConflictMedQA benchmark reveals significant challenges for LLMs in clinical decision-making, primarily their struggle with the temporal dynamics of medical knowledge and internal consistency. Even advanced models, adept at endorsing current guidelines, often faltered markedly when required to reject outdated advice. This asymmetry, coupled with the finding that larger model scale does not consistently reduce internal knowledge conflicts, suggests that unique complexities arise in this domain beyond standard NLP capabilities. These issues, especially prevalent in areas like therapeutic recommendations, could pose direct risks if LLMs are integrated into clinical workflows without a deep understanding of their failure modes.
Investigating mitigation strategies offered further insights. While RAG generally improved adherence to current information, its utility was nuanced. Notably, for smaller models, RAG alone could paradoxically degrade their ability to reject outdated advice, suggesting that merely providing external information can be counterproductive if the model lacks the capacity to critically discern and integrate it, potentially overwhelming weaker internal knowledge structures. This indicates that effective retrieval is as much about the model’s ability to use information as it is about accessing it.
DPO offered a simple complementary approach, demonstrably enhancing alignment with current guidelines and reducing internal conflicts. However, these improvements in complex clinical scenarios stood in contrast to the near-perfect performance models presumably achieve on the specific raw medical advice pairs used during DPO training. This discrepancy suggests a significant challenge in generalizing knowledge learned from such simple pairs to the multifaceted reasoning required in clinical practice, hinting at a gap between memorized correct responses and their robust, contextual application.
The most promising path appears to be the synergistic combination of these approaches. Our findings show that RoD, applying RAG to DPO-tuned models, yielded substantial improvements across all metrics, particularly in enhancing smaller models’ rejection of outdated advice and minimizing internal conflicts. These gains always exceed the sum of the individual contributions from RAG-only or DPO-only applications. While models may struggle to effectively apply DPO-learned parametric knowledge across diverse and complex scenarios, the integration with RAG appears pivotal. Knowledge retrieved via RAG seems to activate relevant DPO-instilled parametric knowledge within the model, leading to these markedly enhanced outcomes and avoiding the potential side effects of RAG-only or the more modest improvements from DPO-only strategies.
These observations also underscore a significant limitation of evaluating LLMs using metrics focused on isolated factual accuracy. The marked performance decline when models face realistic clinical scenarios, which embed cognitive complexities and factors like SDoH, emphasizes the strong need for evaluation methodologies that capture the multifaceted nature of clinical decision-making.
7 Conclusion
Ultimately, for the safe and effective integration of LLMs into clinical practice, future efforts should prioritize the development of robust, hybrid methodologies designed to enhance adaptability to evolving knowledge and ensure internal consistency. This entails creating more contextually rich training and evaluation paradigms that mirror the complexity of real-world clinical encounters, thereby moving beyond isolated assessments to foster genuine contextual understanding and reliability in these critical systems.
Limitations
@@ 7 Conclusion
Ethical Considerations
We have not identified any ethical concerns directly related to this study.
Appendix A Additional Results
A.1 Domain-Specific Models
We further evaluated domain-specific medical models including Med42-8B and Med42-70B Christophe et al. (2024) as well as OpenBioLLM-70B Ankit Pal (2024). Their performance is summarized in Table 2.
Table 2: Results Summary: Domain-Specific Models
Model IKCR ECDA_all ECDA_adh ECDA_rej
Med42-8B (Base) 0.4626 0.5529 0.5497 0.5562
Med42-8B (RAG) 0.4990 0.6914 0.8643 0.5184
Med42-8B (DPO) 0.2056 0.7406 0.6601 0.8210
Med42-8B (RoD) 0.1246 0.8380 0.7902 0.8858
Med42-70B (Base) 0.7055 0.5762 0.8713 0.2811
Med42-70B (RAG) 0.4000 0.7730 0.9566 0.5893
OpenBioLLM-70B (Base) 0.5963 0.5560 0.7298 0.3831
OpenBioLLM-70B (RAG) 0.6053 0.6545 0.9273 0.3818
A.2 LoRA Ablation Studies
We conducted systematic ablation studies to optimize LoRA hyperparameters for DPO fine-tuning, focusing on the rank parameter ($r$) while keeping alpha ($\alpha$) fixed at 16. The results are shown in Table 3.
Table 3: Ablation Results (Mistral-8B)
Rank ($r$) IKCR ECDA_all ECDA_adh ECDA_rej
4 0.2639 0.7524 0.7515 0.7534
8 0.1504 0.8331 0.8145 0.8517
16 0.1181 0.8417 0.7986 0.8848
Overall, higher rank values consistently improve performance given the same training data. Rank 16 achieves the best balance between parameter efficiency and knowledge embedding effectiveness, and the trend suggests that larger ranks enable more effective parametric knowledge injection.
A.3 Factors Impact Analysis
We systematically analyzed how different cognitive factors affect model performance to understand the realistic complexity introduced by our benchmark design. The results are presented in Table 4.
Table 4: Factor-wise Performance Analysis (LLaMA-8B)
Factor Type IKCR ECDA_all ECDA_adh ECDA_rej
Self-Diagnosis 0.4333 0.5462 0.4872 0.6051
Recency Factor 0.4579 0.5462 0.4462 0.6462
Confirmation 0.4833 0.5718 0.5282 0.6154
Frequency 0.4655 0.5308 0.4667 0.5949
Cultural Factor 0.4959 0.5846 0.5487 0.6205
Status Quo 0.4000 0.5308 0.4256 0.6359
False Consensus 0.4273 0.5385 0.4410 0.6359
Racial/Ethnic 0.4915 0.5564 0.5077 0.6051
Socioeconomic 0.4364 0.5256 0.4308 0.6205
Geographic 0.4690 0.5615 0.4872 0.6359
No Factor 0.4444 0.6000 0.5333 0.6667
Models generally achieve higher ECDA scores under the “No Factor” condition, validating our benchmark design. We do not observe systematic bias toward incorrect recommendations despite factor inclusion, indicating that the factors simulate realistic clinical complexity without compromising evaluation validity.
A.4 Comprehensive Performance Visualization
To provide deeper insights into model behavior across different cognitive factors and clinical change types, we present detailed performance breakdowns across all evaluated metrics.
A.4.1 Performance by Cognitive Factor
Figure 5: $ECDA_{adh}$ performance across clinical factors. This metric measures models’ ability to correctly endorse up-to-date medical recommendations under different cognitive biases.
Figure 6: $ECDA_{rej}$ performance across clinical factors. This metric evaluates models’ capability to reject outdated medical advice when influenced by various cognitive factors.
Figure 7: Overall ECDA performance ($ECDA_{all}$) across clinical factors, representing the balanced assessment of both endorsement and rejection capabilities.
Figure 8: Internal Knowledge Conflict Ratio (IKCR) across clinical factors. Lower values indicate better internal consistency, with “No Factor” serving as the baseline condition. The legend below provides symbol/color references.
As shown in Figures 912, the “No Factor” condition consistently yields the best performance across ECDA metrics, aligning with Table 4. Different clinical change types pose varying challenges; in particular, Implementation Approach and Treatment Modality tend to exhibit higher IKCR (cf. Figure 12), indicating greater internal tension for these settings. Larger models do not uniformly outperform smaller ones on rejection ($ECDA_{rej}$; Figure 10), consistent with our hypothesis regarding pre-training bias amplification. Finally, trends in $ECDA_{adh}$ (Figure 9) and $ECDA_{rej}$ (Figure 10) mirror the aggregate $ECDA_{all}$ behavior (Figure 11), supporting the robustness of our evaluation framework.
Figure 9: $ECDA_{adh}$ performance across clinical change types. This metric measures models’ ability to correctly endorse up-to-date medical recommendations under different cognitive biases.
Figure 10: $ECDA_{rej}$ performance across clinical change types. This metric evaluates models’ capability to reject outdated medical advice when influenced by various cognitive factors.
Figure 11: Overall ECDA performance ($ECDA_{all}$) across clinical change types, representing the balanced assessment of both endorsement and rejection capabilities.
Figure 12: Internal Knowledge Conflict Ratio (IKCR) across clinical change types. Lower values indicate better internal consistency, with “No Factor” serving as the baseline condition. The legend below provides symbol/color references.
The tables below detail mitigation effects across different models, clinical factors, and advice change types. Overall mitigation strategy comparisons are presented in Table 5, while Table 6 reports results specific to the confirmation factor.
Table 5: Mitigation Strategy Performance Comparison (Overall)
Model Strategy IKCR ECDA_all ECDA_adh ECDA_rej
Gemma-2-27B Base 0.397 0.580 0.481 0.678
RAG 0.308 (-0.089) 0.759 (+0.179) 0.821 (+0.340) 0.697 (+0.018)
GPT-4o Base 0.612 0.646 0.898 0.395
RAG 0.352 (-0.260) 0.806 (+0.160) 0.965 (+0.067) 0.648 (+0.253)
LLaMA-3.3-70B Base 0.455 0.610 0.662 0.557
RAG 0.297 (-0.158) 0.829 (+0.220) 0.958 (+0.295) 0.701 (+0.144)
LLaMA-3-8B Base 0.456 0.554 0.482 0.626
RAG 0.699 (+0.243) 0.620 (+0.067) 0.935 (+0.453) 0.305 (-0.320)
DPO 0.294 (-0.161) 0.747 (+0.194) 0.772 (+0.290) 0.723 (+0.097)
RoD 0.238 (-0.217) 0.847 (+0.293) 0.934 (+0.452) 0.760 (+0.135)
Mistral-8B Base 0.345 0.553 0.302 0.804
RAG 0.398 (+0.053) 0.744 (+0.191) 0.876 (+0.574) 0.613 (-0.191)
DPO 0.150 (-0.195) 0.833 (+0.280) 0.815 (+0.513) 0.852 (+0.048)
RoD 0.104 (-0.242) 0.888 (+0.335) 0.879 (+0.578) 0.897 (+0.094)
Qwen2.5-7B Base 0.651 0.550 0.745 0.354
RAG 0.509 (-0.142) 0.720 (+0.170) 0.935 (+0.190) 0.504 (+0.150)
DPO 0.437 (-0.214) 0.680 (+0.130) 0.814 (+0.068) 0.546 (+0.192)
RoD 0.263 (-0.388) 0.807 (+0.257) 0.880 (+0.135) 0.734 (+0.380)
Qwen2.5-72B Base 0.710 0.597 0.916 0.278
RAG 0.729 (+0.019) 0.624 (+0.027) 0.982 (+0.066) 0.266 (-0.012)
Table 6: Mitigation Performance for Confirmation Factor
Model Strategy IKCR ECDA_all ECDA_adh ECDA_rej
Gemma-2-27B Base 0.393 0.587 0.523 0.651
RAG 0.302 (-0.091) 0.756 (+0.169) 0.831 (+0.308) 0.682 (+0.031)
GPT-4o Base 0.622 0.626 0.908 0.344
RAG 0.354 (-0.268) 0.818 (+0.192) 0.995 (+0.087) 0.641 (+0.297)
LLaMA-3.3-70B Base 0.465 0.608 0.697 0.518
RAG 0.295 (-0.170) 0.830 (+0.223) 0.964 (+0.266) 0.697 (+0.179)
LLaMA-3-8B Base 0.483 0.572 0.528 0.615
RAG 0.696 (+0.213) 0.623 (+0.051) 0.954 (+0.426) 0.292 (-0.323)
DPO 0.323 (-0.161) 0.754 (+0.182) 0.790 (+0.261) 0.718 (+0.103)
RoD 0.242 (-0.241) 0.846 (+0.274) 0.939 (+0.410) 0.754 (+0.138)
Mistral-8B Base 0.317 0.572 0.339 0.805
RAG 0.397 (+0.080) 0.756 (+0.185) 0.897 (+0.559) 0.615 (-0.190)
DPO 0.153 (-0.163) 0.839 (+0.267) 0.821 (+0.482) 0.856 (+0.051)
RoD 0.099 (-0.218) 0.892 (+0.321) 0.877 (+0.538) 0.908 (+0.103)
Qwen2.5-7B Base 0.671 0.541 0.744 0.339
RAG 0.535 (-0.136) 0.715 (+0.174) 0.944 (+0.200) 0.487 (+0.149)
DPO 0.456 (-0.215) 0.659 (+0.118) 0.790 (+0.046) 0.528 (+0.190)
RoD 0.263 (-0.408) 0.821 (+0.279) 0.887 (+0.144) 0.754 (+0.415)
Qwen2.5-72B Base 0.742 0.595 0.944 0.246
RAG 0.732 (-0.010) 0.633 (+0.038) 0.995 (+0.051) 0.272 (+0.026)
A.5 Recommendation Intensity Category: Clinical Justification
Addressing concerns about the clinical validity of recommendation intensity modifications, we provide detailed justification for this category’s inclusion and its impact on our benchmark.
Clinical Significance of Intensity Variations. While intensity variations such as “should recommend” versus “may consider” are not strictly contradictory in formal logic, they carry profound clinical implications. First, clinical studies demonstrate that “should” language typically results in adherence rates of approximately 80%, compared to only 20% when phrased as “may consider.” Second, many real-world clinical guideline updates explicitly focus on the strength of recommendation rather than altering the core intervention. Finally, practice variation studies show that intensity changes directly influence clinical decision-making patterns and patient outcomes.
Example Analysis. For instance, a current recommendation such as “People without immunity should receive full vaccination” differs substantially in clinical impact from a modified version: “People without immunity may consider receiving full vaccination.” This shift constitutes a meaningful clinical conflict that affects patient outcomes and public health recommendations, and accounts for 27.2% of our dataset scenarios.
A.6 External vs. Internal Conflict Framework
To clarify our conflict detection methodology, we distinguish between external and internal conflicts.
External Conflicts. Each recommendation pair ($R_{\text{current}}$, $R_{\text{outdated}}$) generates scenarios $S_{\text{current}}$ and $S_{\text{outdated}}$, which are evaluated independently against current medical ground truth. An external conflict occurs when the model endorses $S_{\text{outdated}}$ (which should be rejected) or rejects $S_{\text{current}}$ (which should be endorsed).
Internal Conflicts. These are assessed using paired scenarios where simultaneous endorsement indicates internal knowledge inconsistency. For a given scenario pair ($S_{i,\text{current}}$, $S_{i,\text{outdated}}$) for concept $i$, an internal conflict arises when the model endorses both scenarios. The Internal Knowledge Conflict Rate (IKCR) quantifies the frequency of such contradictions across all active pairs, as reported in Table 2.
A.7 Analysis of Counterintuitive Scale Effects
Our investigation revealed unexpected patterns where larger models sometimes underperform smaller variants, particularly in rejection tasks.
Empirical Evidence. Table 7 shows representative results across three model families, highlighting that parameter scaling does not guarantee improved performance on ECDA_rej.
Table 7: Scale Effects on ECDA_rej Performance
Model Family Parameter Size ECDA_rej Performance
Qwen 7B $\to$ 72B 0.3540 $\to$ 0.2783
LLaMA 8B $\to$ 70B 0.6256 $\to$ 0.5571
Med42 8B $\to$ 70B 0.5562 $\to$ 0.2811
Proposed Mechanistic Explanation. We hypothesize this phenomenon results from pre-training bias amplification. Clinical scenarios rich in specialized terminology may trigger strong correctness associations learned during pre-training (the authority signal hypothesis). Larger models, exposed to broader corpora, develop stronger heuristic associations between clinical language and authoritative content, leading to scale-dependent bias. These pre-training biases can override rejection capabilities acquired during RLHF or instruction tuning, especially when plausible but incorrect recommendations are presented. By contrast, smaller models may be less affected due to weaker initial biases and a proportionally greater influence of alignment training updates. This observation emphasizes that medical LLM evaluation requires careful consideration of both capability scaling and bias amplification effects.
Appendix B Detailed Description of LLMs
Below we provide a brief description of each large language model (LLM) evaluated in our study, highlighting their key architectural and training characteristics.
GPT-4o is OpenAI’s multimodal model. While the exact parameter count remains undisclosed, GPT-4o features a unified architecture capable of processing and generating text, images, and audio with a context window of up to 128,000 tokens. It achieves comparable or better text performance relative to GPT-4, but with significantly lower latency and cost. The model is instruction-tuned and optimized for real-time interactive applications. We used GPT-4 via the OpenAI API under its terms of use.
Llama-3-8B and Llama-3-70B are Meta’s latest open-weight models, featuring 8 billion and 70 billion parameters, respectively. Both are dense decoder-only Transformers trained on approximately 15 trillion tokens of deduplicated public data. Instruction-tuned versions incorporate multi-stage reinforcement learning from human feedback (RLHF), and Meta provides both default (8K) and long-context (up to 128K) variants for research.
Qwen2.5-7B and Qwen2.5-72B are Alibaba’s state-of-the-art models with 7 billion and 72 billion parameters. Qwen 2.5 introduces a greatly expanded pre-training corpus (18T tokens) and large-scale supervised fine-tuning (over 1 million samples), along with reinforcement learning and reward modeling. Both models natively support a 32,000-token context window.
Gemma-2-27B-it is Google DeepMind’s 27-billion-parameter, instruction-tuned model from the Gemma 2 family. It employs dense Transformer architecture with interleaved local-global attention and group-query attention to improve memory efficiency. Gemma-2 models are trained on up to 8T tokens and are designed for efficient inference on single high-memory GPUs or TPUs, released under the Apache 2.0 license.
Ministral-8B-Instruct-2410 is a recently released model from Mistral AI, designed for local and on-device use. It features 8 billion parameters with a dense Transformer architecture and a context window of up to 128,000 tokens, enabled by interleaved sliding-window attention. Ministral-8B-Instruct
Appendix C More details in Dataset Construction